OpenStudy (anonymous):

WILL GIVE A MEDAL PLEASE HELP What is the equation of a line with a point (-3, 0) and no slope?

OpenStudy (anonymous):

-3+0

OpenStudy (anonymous):

Answer choices: A. x + 3 = 0 B. y + 3 = 0 C. x - 3 = 0 D. y - 3 = 0

OpenStudy (anonymous):

A